OpenStudy (anonymous):

An electron and a proton have charges of an equal magnitude but opposite sign of 1.60 10-19 C. If the electron and proton in a hydrogen atom are separated by a distance of 1.50 10-11 m, what are the magnitude and direction of the electrostatic force exerted on the electron by the proton?

OpenStudy (irishboy123):

Coulombs law: F = k q1 q2 / r^2 k = 9 x 10^9 ok?

OpenStudy (ybarrap):

Remember also that positive charges repel each other and negative charges repel, as well. So if in @IrishBoy123's equation above you get a positive number for the force, that means that the forces act to "push" the charges away from each other. Conversely, if F turns out to be negative, because one is positive and the other is negative, then the charges will attract each other. That is the significance of the sign (plus or minus) in his equation.

OpenStudy (shamim):

Given q1=q2=1.6*10^-19C r=1.5*10^-11m F=?

OpenStudy (shamim):

Use coulomb's law

OpenStudy (shamim):

Response plz!!!!!!!

OpenStudy (anonymous):

@ybarrap Great explanation on when the forces attract or repel! It helped a lot. Thank you! Also thank you @IrishBoy123

OpenStudy (ybarrap):

You're welcome